OpenStudy (anonymous):

3 is a root of the equation x^2-9x+c=0 a) 12 b) 6 c) -6 d) 5

OpenStudy (lgbasallote):

so what's the question? what's c?

OpenStudy (lgbasallote):

i can solve this using remainder theorem...or by division of polynomials. pick your poison

OpenStudy (anonymous):

i choose the polynomials thingy..lay it on me

OpenStudy (lgbasallote):

if 3 is a root that means x -3 is a factor so divide ______________ x-3 | x^2 - 9x + c the remainder here should be 0..anyway moving on... x - 6 ______________ x-3 | x^2 - 9x + c x^2 - 3x ----------- -6x + c -6x + 18 like i said the remainder should be 0 so c - 18 = 0 solve for c c = 18 that should be the answer..are you sure about the question and the chocies?

OpenStudy (lgbasallote):

i can prove my answer by remainder theorem x - 3 is a factor so p(3) = 0 p(3) = (3)^2 - 9(3) + c = 0 9 - 27 + c = 0 -18 + c = 0 c = 18

OpenStudy (lgbasallote):

soa gain are you sure it's looking for c and those are the choices>

OpenStudy (lgbasallote):

so again*

OpenStudy (anonymous):

oh yeah the question is find the other root so i got stuck with understanding wat to do with c

OpenStudy (lgbasallote):

or it could be looking for the other root

OpenStudy (lgbasallote):

ahh yes..i was right

OpenStudy (lgbasallote):

so c = 18 x^2 - 9x + 18 = 0 one of the factors is x - 3 so divide \[\frac{x^2 - 9x + 18}{x-3} \implies \frac{(x-3)(x-6)}{x-3}\] do you see it?

OpenStudy (anonymous):

yes i think i kinda get the picture i just need to load it

OpenStudy (lgbasallote):

haha go for it! once you get the quotient just solve for x

OpenStudy (lgbasallote):

then you have the answer. i have to go now so sorry i cannot explain further

OpenStudy (lgbasallote):

but here's an example to guide you \[\frac{(x+2)(x-1)}{x-1} \implies x + 2\] solve for x x + 2 = 0 x = -2

OpenStudy (anonymous):

can i just ask the questions ? is sensible to replace 0 with 3?

OpenStudy (anonymous):

hello @lgbasallote ?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

like this x^2-9x+c=3

OpenStudy (lgbasallote):

no it's not sensible

OpenStudy (lgbasallote):

if it's x^2 - 9x + c -3 = 0 then it's sensible to do that

OpenStudy (lgbasallote):

okay i really have to go. good luck!

OpenStudy (anonymous):

i need it